You want to purchase the French Government Bond OAT. You are looking for a maturity between 5-10 years. The initial margin requirement in this maturity range is 10% of market value. The bond you select is trading at 115.00 with a face value of EUR 1000. You wish to invest EUR 1000. How many bonds can you purchase, given the margin requirement?


a)5

b)8

c)9

d)10



You want to purchase the French Government Bond OAT. You are looking for a maturity between 5-10 years. The initial margin requirement in this maturity range is 10% of market value. The bond you select is trading at 115.00 with a face value of EUR 1000. You wish to invest EUR 1000. What is your resulting total exposure and actual margin requirement?


a)EUR 10,000 for a margin of EUR 1000

b)EUR 8,000 for a margin of EUR 800

c)EUR 9,200 for a margin of EUR 920


Thank you :)


1
Expert's answer
2021-02-17T14:51:48-0500

1. The maximum amount to be purchased with 1,000 is 115×8 = 920.

So, the correct answer is b) 8.

The resulting total exposure and actual margin requirement is:

115×8 = 920.

920/0.1 = 9,200.

So, the correct answer is c) EUR 9,200 for a margin of EUR 920.